OpenStudy (anonymous):

cos20.cos40.cos60.cos80 = 1/16

OpenStudy (anonymous):

cos20.cos40.cos60.cos80 = 1/16

OpenStudy (anonymous):

20,40,80 is 1/8 ( http://en.wikipedia.org/wiki/Morrie%27s_law) so 1/8 *1/2 = 1/16

OpenStudy (anonymous):

cos100 + cos 20

OpenStudy (anonymous):

(sinasin2a+sin3asin6a+sin4asin13a)/(sinacos2a+sin3asin6a+sin4acos13a)= tan9a